You invest $30,000 in an account for your retirement. the interest you receive is 8% compounded monthly.
A. How much will be in the account after 30 years?
B. If your interest was compounded continously at the same rate, how much would you have after 30 years?
PLease help!

Answer by stanbon(48558)

You invest $30,000 in an account for your retirement. the interest you receive is 8% compounded monthly.
A. How much will be in the account after 30 years?
A(30) = 30000(1+0.08/12)^(12*30)
A(30) = 30000(1.00666666...)^360
A(30) = 30000(10.93572966...)
A(30) = $328,071.89
-----------------
B. If your interest was compounded continously at the same rate, how much would you have after 30 years?
A(30)=30000e^(0.08*30)
A(30)=30000(11.02317638...)
A(30)=$330,695.29
--------------------
Cheers,
Stan H.